OpenStudy (anonymous):

** i give medals** Which of the following ordered pairs is the solution to the system of linear equations? x plus y equals fourteen, Two x plus y equals twenty-four (7, 7) (3, 11) (6, 12) (10, 4)

OpenStudy (anonymous):

10+4=14 2*10+4=24 (10,4)

OpenStudy (cwrw238):

x + y = 14 2x + y = 24 take each ordered pair and see which one fits

OpenStudy (igreen):

x + y = 14 2x + y = 24 Let's plug in (10, 4) into both equations and see if it's correct: x + y = 14 10 + 4 = 14 14 = 14 So that fits in the first equation, now let's try the 2nd: 2x + y = 24 2(10) + 4 = 24 20 + 4 = 24 Can you add 20 and 4? @kyrea

OpenStudy (cwrw238):

e eg (7,7) 7 + 7 = 14 2(7) + 7 = 21 so its not this one

OpenStudy (igreen):

@roz_mar23 Don't give out answers.

OpenStudy (anonymous):

its 6/12

OpenStudy (anonymous):

@linaboo123 no because 6+12 doesnt =14

OpenStudy (anonymous):

my teacher gave me that answer

OpenStudy (anonymous):

so which one is it

OpenStudy (anonymous):

its 6/12 my teacher gave me the answer

OpenStudy (igreen):

Add 20 and 4. @kyrea

OpenStudy (igreen):

(6, 12) is WRONG, @linaboo123

OpenStudy (igreen):

What's 20 + 4? @kyrea

OpenStudy (anonymous):

THIS IS HER MATH TEACHER BECKY 6/12 IS THE ANSWER

OpenStudy (anonymous):

lol @iGreen i understand u want me to work it out but i just wanted the answer

OpenStudy (igreen):

@kyrea We're not here to give answers, we're here to guide you. If you chose @linaboo123 's answer, it was wrong, plus she didn't follow the Code of Conduct.

OpenStudy (anonymous):

yea ok whatever 👋